Bramble Corp. sells two types of computer hard drives. The sales mix is 30% (Q-Drive) and 70% (Q-Drive Plus). Q-Drive has variable costs per unit of $60and a selling price of $120. Q-Drive Plus has variable costs per unit of $75 and a selling price of $165. Bramble’s fixed costs are $1134000. How many units of Q-Drive would be sold at the break-even point?

a)14000.

b)4200.

c)5670.

d)9800.
Answer #1

Combined contribution margin

= [(120-60)*30%] + [(165-75)*70%]

= 18 + 63

= 81

Breakeven point = Fixed cost / combined contribution margin

= 1134000/81 = 14,000

Q-Drive units in break even point = 14,000*30%

= 4,200

Option B